### v3

Abstract
Conformal phased arrays provide multiple degrees of freedom to the scan angle, which is typically limited by antenna aperture in rigid arrays. Silicon-based RF signal processing offers reliable, re-configurable, multi-functional, and compact control for conformal phased arrays that can be used for on-the-move communication. While the lightweight, compactness, and shape-changing properties of the conformal phased arrays are attractive, these features result in dynamic deformation of the array during motion, leading to significant dynamic beam pointing errors. We propose a silicon-based, compact, reconfigurable solution to self-correct these dynamic deformation-induced beam pointing errors. Furthermore, additive printing is leveraged to enhance the flexibility of the conformal phased arrays, as the printed conductive ink is more flexible than bulk copper and can be easily deposited on flexible sheets using different printing tools, providing an environmentally-friendly solution for large-scale production. The inks such as conventional silver inks are expensive and copper-based printable inks suffer from spontaneous metal oxidation that alters trace impedance and degrades beamforming performance. This work uses a low-cost molecular copper decomposition ink with reliable RF properties at different temperature and strain to print the proposed intelligent conformal phased array operating at 2.1 GHz. Proof-of-concept prototype $2\times 2$ array self-corrects the deformation induced beampointing error with an error $<1.25^{\circ}$. The silicon-based array processing part occupies only 2.58 mm2 area and 83 mW power per tile.
keywords:
Additive Printing, conformal phased arrays, deformation correction, beam stabilization, molecular copper decomposition ink, self-correction loop, scalable tile-based.
1 Introduction
Phased array systems support directional data transfer, which enables inherent spatial selection to support ever-increasing wireless data rate. Initial research on the phased arrays started several decades back (earliest reported in 1909) to enable inertia-free fast beam steering by avoiding mechanical gimballing [1, 2]. The underlying assumption in these works was that the antenna surfaces were rigid. By the 1940s, conformal arrays were developed to curve around spherical and nonuniform surfaces[3, 4, 5, 6] that afforded array systems an added degree of freedom by eliminating the conventional radome. Conformal phased arrays have been used extensively since then in the automotive, aviation, and space industries. Having a non-uniform antenna structure enhances the field of view (FoV). When deployed on an aircraft, the array benefits from the aerodynamic shape of the plane, which improves the fidelity of the in-flight wireless and makes the airborne systems such as Unmanned Aerial Vehicles (UAV) management easy, as shown in Fig. 1.
Satellite communication (SATCOM) has embraced conformal phased arrays due to their lightweight, small volume, and shape-changing properties. The manageable costs associated with launching and deploying [7, 8] have led to the widespread adoption of conformal phased arrays in Low-Earth Orbit (LEO) satellites[9]. LEO satellite networks such as Amazon Kuiper[10], SpaceX Starlink [11], OneWeb, and Telesat Light-speed, boast coverage fields that collectively span the entire terrestrial surface area, opening profound opportunities for direct-to-consumer wireless communications in remote areas [12, 13]. These satellite networks, in conjunction with wearable/textile arrays[14, 15, 16], will further aid challenging research expeditions and facilitate disaster relief management as illustrated in Fig. 1.
Figure 1: Dynamic wing load variations and wing vibrations lead to deformed antenna arrays impacting in-flight wireless and navigation in Unmanned Aerial Vehicles (UAV); dynamic deformation in SATCOM and lightweight textile arrays. Silicon-based self-correction processor addresses gain loss and beam pointing error due to dynamic deformation in conformal phased arrays in real-time with the proposed additively printed, lightweight, tile-based flexible array.
Surface deformation of a conformal phased array with curvature enhances its Field-of-View (FoV) but incurs increased inter-element coupling [17] and path length variations leading to beam pointing errors [18]. For satellites, it is possible to estimate the amount of deformation following the expansion of arrays once they are placed in orbit. While the initial deformation due to the aerodynamic shape of aircraft and drones can be estimated, the dynamic deformations arising from wing loading and vibrations[19] during transit cannot be predetermined. These dynamic deformations depend on factors such as array weight, size, drone/aircraft speed, wind conditions, initial deformation, etc. Intelligent self-adaptive conformal phased arrays will serve as the cornerstone to enable existing and emerging applications, as shown in Fig. 1, while overcoming limitations from dynamic deformations [20].
Recent studies have explored numerous strategies for correcting the radiation pattern of conformal phased arrays [21, 22, 23] arising from physical deformities. However, element-level deformation sensing and compensation with conventional mechanical strain sensors [24, 25] and sensing mutual coupling [26, 27] has been ineffective due to complexity and non-planar shapes beyond single-point curvature. Machine-learning (ML), and in particular, pre-trained deep-learning-based networks, are increasingly being used for adaptive beam synthesis, allowing arbitrary shape correction based on training[19]. However, it relies heavily on feature-rich longitudinal-series training data with prior knowledge of the deformation surface and uncertain operational environments. These models further require separate dictionary sets for near-field and far-field patterns as well as environmental conditions, demands exponential increase in storage as the array size is scaled. The multi-dimensional variable dynamic range search algorithm in [28], employs array-level compensation suited for both near- and far-field. However, its higher computational complexity and the additional requirement for separate generation and recovery units with high computing power along with low-power transceivers make this technique challenging to adapt for airborne platoons and portable units. Phase correction using iterative methods such as genetic algorithm [29] and iterative phase synthesis [30, 31, 32] not only requires external engines for executing the iterative algorithm, but also incurs large latencies not suited for fast communications-on-the-move applications.
In this article, we present a self-adaptive conformal phased array additively printed using a novel copper molecular decomposition ink with inherent deformation correction from both material (additively printing-based) and physical deformation effects. While the ink shows outstanding electrical performance capabilities matching the performance of the traditional chemical etching process, and exhibiting stable RF characteristics under varied temperatures and strain conditions, the silicon-based calibration technique effectively compensates for dynamic deformities, irrespective of the nature of deformation.
Figure 2: Tile based array scaling with edge computing for gain, phase, delay and impedance self-calibration loops (conceptual). The proposed additively printed $2\times 2$ conformal tile layer stackup with beamforming IC and a. Full break down of Inks in each layers of array, Ball Milled Cu Ink b. SEM (Scanning Electron Microscope) of the ink before sintering, and after sintering c. Test setup for resistance measurements at different temperatures: Printed traces using the copper molecular decomposition ink are placed on a hot plate, and the temperature is varied to note the variation in resistance with respect to temperature. d. Measured resistance at different temperatures.
2 Results and discussions
2.1 Low-cost reliable molecular copper decomposition ink for additive printing
Additively printed RF circuits have gained prominence in next-generation wireless systems with their ability to generate non-uniform conformal shapes required for large-scale infrastructure while being environmentally-friendly [33, 34, 35, 14, 15, 16]. Recent works for additive printing techniques such as aerosol jetting, ink jetting, direct writing, and screen printing have shown enhanced scalability, reduced material waste generation, and cost-effective manufacturing processes. However, the efficacy of the additive printing method is heavily contingent upon the quality and properties of the conductive inks utilized in the application. Common challenges in additive manufacturing of conductive inks revolve around printability and electrical performance [36]. Factors such as additives, ink viscosity, uniformity, micro/nanomaterial structure, and size directly influence the potential printing method, substrate material, sintering conditions, and electrical performance of the traces. The electrical performance of printed conductive traces is also commonly evaluated for stability under repeated exposure to high temperatures, corrosive environments, and mechanical bending and torsion, reflecting long-term reliability during use.
Current developments of the ink drives a cheaper and higher performance alternative to existing metallic, polymer, and carbon-based inks that are increasingly costly or fail to meet the electrical performance standards achieved through conventional manufacturing methods [37]. When considering conductive inks, metallic-based inks consistently demonstrate superior performance in RF electronics due to their inherently higher conductivity, thermal conductivity, and suitability for various additive manufacturing methods. Metallic-based inks typically fall into two categories: metallic nanostructured-based inks and molecular precursor-based inks. The former often involve nanoparticles, nanowires, or nanoplates as the fundamental conductive element, while the latter are typically in a precursor form that is reduced during sintering to form the metallic state of the salt.
Figure 3: a. Dipole antenna printed using proposed Cu decomposition ink testing under varied strains b. Laser ablated dipole antenna tested at different temperatures on a hotplate c. Test set up for measuring ink stability under different strains. d. Measured S11 of the dipole antenna printed using Cu ink to demonstrate ink stability at different temperatures and e. strain.
Although silver (Ag)-based inks have been largely appreciated due to their low resistivity (1.59 $\mu\Omega.cm$) for use in printed electronics, their high-cost margins make copper (Cu) a much more cost-effective alternative with potential for comparable resistivity performance (1.72 $\mu\Omega.cm$) [38, 39]. However, excessive additives needed to improve printability of Cu nanostructured inks and with a greater potential for ambient condition oxidation suggests that molecular Cu-based inks are more advantageious for long term commercial and industrial applications [39, 40, 41]. Herein a molecular Cu formate-based ink is utilized to create a highly conductive Cu thin-film ($=$35 MS/m)[40] on a Pyralux substrate with demonstratively consistent electrical, and RF performance under varied strains and temperatures.
Figure 4: a. 4-element beamforming receiver architecture with four downconverters and discrete-time beamformer with self-corrected calibration loop in feedback. The beamformer output is digitized using an off-chip ADC and fed back to the on-chip calibration loop. b.A single tile self-calibration loop structure for correcting beam pointing error c. Single channel receiver architecture with LNTA, downconversion mixer, phase shifter, TIA, and time-delayed sampler. d.Illustration of self-calibration loop finding optimum phase control code under dynamic deformation.
2.2 Scalable tile-based delay-phase tunable receiver array
The emergence of highly integrated silicon-based RF signal processors has revolutionized array synthesis and beam steering with multi-modal spatial signal processing [42, 43, 44], enabling high Signal-to-Noise Ratio (SNR) in commercial arrays. When the number of elements increases, the beam width reduces, thereby enhancing the spatial selectivity. At 2.1 GHz, the antenna dimensions are in centimeters, and therefore, scaling to a larger array is not straightforward. This scaling comes with several challenges [45], such as additive printing of high-resolution structures with physically large dimensions and combining signals from multiple antennas over wide bandwidth with high signal-to-noise ratio. We are adopting a modular tile-based approach for the receiver array as shown conceptually in the illustration in Fig. 2. In this work, each tile consists of a $2\times 2$ conformal antenna array attached to a 4-channel beamforming integrated circuit (BFIC). Figure 5e shows four tiles thermally bonded together to form a 16-element array. The BFIC in each tile combines each antenna signal and, therefore, acts as a sub-array. Each sub-array outputs are then combined using a second-stage combiner to achieve $4\times 4$ radiation pattern. The gain, phase shift, and impedance mismatch, and delay skews exacerbates as the array scales. The modular tile-based array structure thus enables easier calibration of larger arrays. An N-element array can be divided and processed into N/4 sub-arrays.
The BFIC in a $2\times 2$ tile performs discrete-time true-time-delay BFIC as shown in Fig. 4a. The developed ink and the printing technique ensure that the four channels of the BFIC are connected to each antenna through impedance-matched via holes and traces. Each receiver front end provides a low-noise impedance matching to the antenna/PCB trace impedance. The receiver channels are capable of supporting higher modulation schemes, including M-QAM (Quadrature Amplitude Modulation) and OFDM (Orthogonal Frequency-Division Multiple Access). To process the higher modulated signals, the receivers down-convert the incoming RF signals to the baseband while extracting the in-phase (I) and quadrature-phase (Q) components. Beamforming of the four channels is performed in the baseband after I/Q extraction using a discrete-time beamformer [46]. Time-delayed sampling of the extracted I/Q baseband signals, along with the LO (Local Oscillator) phase shifter, captures coherent samples from each element and enables wideband beamforming as shown in Fig. 4a [47]. Independent gain, phase, and delay tuning is provided for each of the I and Q channels to reduce the respective mismatches in each channel. A high-frequency external clock (2×2$\times$2 ×RF input frequency) is provided for the signal down-conversion and time-delayed sampling, which is susceptible to impedance variations due to deformation and printing imperfections. A tunable impedance matching network is used at the clock input to alleviate these impedance imperfections.
The PCB stack-up for the prototype conformal tile is shown in Fig. 2, with the 4-channel BFIC attached to a $2\times 2$ additively printed array with the backplane antenna. The additively printed conformal array comprises of flexible sheets of multiple DuPont Pyralux® AP (polyimide) and Ninjaflex substrates. The BFIC and its RF and non-RF traces are printed over AP layers of thickness 0.127 mm, combined using 0.0508 mm Fast Rise EZ (epoxy). The antenna ground plane is printed over 0.254 mm AP combined with other layers using 0.0508 mm Fast Rise EZ. The antennas have been printed on a Ninjaflex substrate with 7.62 mm thickness, chosen to meet the ground plane height requirements at 2.1 GHz. Circular patch antennas are used, as bending has minimal impact on the return loss and gain. The three DuPont Pyralux® AP sheets are combined using Fast Rise EZ as mentioned, which are then thermally bonded with the Ninjaflex substrate.
A full breakdown of inks used in each layer are also shown in Fig. 2. The implemented BFIC is on layer 5 with Cu ink printed non-RF traces on layer 4. Cu clad is used for RF traces on layer 3 and antenna reflector on layer 2. Finally, Ag (silver) ink is used to print antennas on layer 1 (shown in Fig. 6e). Ag vias and Cu Ink vias that were fixed with Ag are used between the layers.
2.3 On-chip dynamic deformation correction
Beam pointing errors due to deformation, as explained in Section 1, can lead to loss in communication, even a small misalignment can drastically degrade the SNR. We propose a low-power, low-area silicon-based self-calibration technique inspired by model-free techniques such as perturb and observe and extremum-seeking[48, 49] principles.
The BFIC output is fed into an integrated self-calibrating loop considering the fact that a beamformer has only one maximum point, which corresponds to the main lobe. The side lobes are considered as local maxima. For each angle, corresponding to each deformation condition (single- or multiple-curvature), there is a unique phase shift combination (and corresponding phase code word) for each element. For an 8-bit phase shifter, there are thus $2^{8}$ possible phase shifts. At a given AoA and a given deformation (single- or multiple-curvature), for a $2\times 2$ array, there are $2^{8\times 3}$ phase-shifting options, where one element is taken as reference (labeled as REF receiver in Fig. 4a). Only one combination among these gives the maximum SNR when the beam is perfectly aligned with the transmitter. Incremental time-based search counters to validate each option is impractical for aforementioned applications-on-the-move. A data-driven or model-driven ML model can provide the optimum combination quickly but gets easily stuck into an unknown condition considering the plethora of uncertainties in real applications. The proposed integrated loop ensures fast convergence to the optimum phase shifter combination.
Figure 4b show the self-calibration loop. The deformation correction process is captured in the following steps illustrated in Fig. 4d:
  • Step 1: Sinusoidal perturbations 1 are generated by the global on-chip look-up table (LUT).
[…]
  • Step 4: The high-pass filter (HPF) and the multiplier 3 estimate the slope and the following accumulator estimates the subsequent step size dynamically 4. The dynamic step size generation ensures fast convergence.
  • Step 5: After recursive iterations, each receiver settles with the phase shifter combination that corresponds to maximum SNR.
The following part of this section explains the proposed loop operation for determining the optimal phase control code $P_{\phi\text{,opt}}$. For simplicity we consider only 2 -elements, considering first element as reference (REF). The phase shifter of the second element is initialized with $P_{\phi\text{,int}}$. The LUT generated perturbation is added to the initial phase control code ($P_{\phi,int}+a_{\phi}sin[\omega\text{n}]$ ) resulting in a sinusoidal change in phase as shown in Fig. 4d. The perturbation is amplified by $a_{\phi}$ based on the convergence requirement when adding multiple calibration loops. Subsequently, the resulting $BF_{\text{out}}$ is applied to the HPF to center it around zero mean, facilitating the capture of resultant variations in $BF_{\text{out}}$. The filter cutoff frequency is deliberately set to a low value to ensure the capture of slow-varying perturbations. The captured $BF_{\text{out}}$ variations are then multiplied with initial perturbations to discern the slope. A correction factor $\Psi$ is needed to compensate the test bench delay, so that both signals will be in phase at the multiplier input. The sign and amplitude of this multiplier output determine the direction and perturbation amplitude (step size) for the subsequent set of iterations (more detailed calculations are shown in Appendix B). The multiplier output is then cumulatively added with the past iteration values to determine a new step size, which is then incorporated into the phase control code, as depicted in Fig. 4d. As the phase control code approaches $P_{\phi\text{,opt}}$, the change in $BF_{\text{out}}$, and consequently, the multiplier output diminishes, signifying that the loop is nearing convergence to $P_{\phi\text{,opt}}$. Figure 4d illustrates convergence to $P_{\phi\text{,opt}}$ in both $P_{\phi\text{,opt}}>$ $P_{\phi\text{,int}}$ and $P_{\phi\text{,opt}}<$$P_{\phi\text{,int}}$ scenarios. It is noteworthy that any dynamic changes in deformation may shift the optimal phase control code, prompting the loop to adapt and self-adjust the phase control code accordingly.
3 Methods
3.1 Ink design details
The copper thin-film material was made from a copper formate-based ink slurry developed through ball milling. The ink included a ratio of 0.45:5:0.45 of copper formate, di-ethylene glycol butyl ether (DEGBE), and dimethylformamide (DMF) respectively in a ceramic ball milling container at an rpm of 300 for 1 hour. The respective slurry was collected in a centrifuge tube after the mechano-chemical synthesis and centrifuged at 6000 rpm for 5 minutes and decanted to remove excess solvents. It was consequently redispersed in 2 mL of DEGBE and centrifuged and decanted once more to remove any excess DMF in solution. Fig. 2b shows the ink before sintering, and after sintering. Prior to sintering the ink is made of an approximate Cu formate flake size distribution from 1-25 µm. After sintering, the copper formate flakes reduce and form a densely percolated thin film of copper nanoparticles.
Figure 2c shows the test setup for validating the electrical properties of the ink at different temperatures. The ink remains stable and shows $<3\%$ change in resistivity across 50 C temperature change. Fig. 3a shows the test setup for validating the stability and RF properties of the ink under various strains. A simple stretchable dipole antenna is printed using the proposed ink and captured the S11 under various strains from 4mm to 9mm. Similarly, the S11 captured at various temperature from the test setup shown in Fig. 3b. The RF properties and ink stability at the frequency of interest (2.1 GHz) is verified under various temperature and strain as shown in Fig. 3d and Fig. 3e.
3.2 IC design details
Both BFIC and deformation correction loop are designed and fabricated in TSMC 65nm GP CMOS process. Direct conversion (homodyne) architecture is adopted for each receiver channel in the BFIC, thus down-converting the RF signal directly to the baseband. As shown in Figure 4c, the front-end of each channel is an inductively degenerated low-noise transconductance amplifier (LNTA) for impedance matching with the antenna. The LNTA is followed by a double balanced current-mode passive mixer driven by four-phase LO clocks. The double balanced structure of mixer enables I/Q extraction from the incoming signal as well as rejects the LO feed through in both RF and IF port. This helps to avoid the offset error in the baseband signal due to self-mixing as well as LO induced baseband distortion. The down-converted quadrature outputs are then applied to a tunable trans-impedance amplifier (TIA) with 3-bit gain calibration( 7dB to 10dB as shown in Fig. 6e) . The TIA is designed using a single-stage pseudo-differential amplifier stabilized using a common-mode feedback circuit. Further, a 3-bit tunable capacitor bank provides first-order channel (blocker) filtering up to 70 MHz for each of quadrature paths separately. A second-order blocker filtering is provided by the large capacitance at the input of the TIA which bypasses the blocker component from the baseband signals to ground. A p-channel source-follower buffer (BUF) followed by the TIA is used to drive baseband discrete-time beamformer.
As mentioned in Section 2.2, the beamformer is realized using a discrete-time sample-and-hold architecture. The beamforming network consists of time-interleaved switched-capacitor array that samples signals from each channel. Based on the AoA and the delay experienced by each element, the sampling instance can be adjusted to extract coherent samples from each channel. This time delayed sampling is enabled by the phase-shifter (PS) in the LO path with the phase-interpolator (PI) in the sample-and-hold clock path which provides true-time-delay and enables beamforming over a wide fractional bandwidth without beam-squint issue[47]. Th 8-bit PS, designed using an inverter-based vector modulator architecture, offers $5.6^{\circ}$ phase resolution in each of the four quadrants. The inverter-based PI in the sampler clock provides 6-bit tunability with a delay resolution of 76 ps. This work considers deformation correction over $\approx$70 MHz base-band bandwidth which requires the PS tuning only for correcting the beam-pointing error. Both PI and PS are designed in the clock path away from the signal path, to minimize direct noise injection into the signal.
The captured coherent samples are integrated over a clock period for beamforming. A two-stage internally compensated OTA with 685MHz unity-gain bandwidth is designed for the beamforming. A 4.2 GHz external clock feeds the on-chip LO and sampler clock generator with a 4-bit tunable matching network as mentioned in Section 2.2. The clock generator includes static clock dividers, clock drivers, and regenerator circuits to minimize power consumption. The prototype BFIC occupies 2.56 $\text{mm}^{2}$ (Fig. 6a) area (including pads) and consumes 18.5 mW per channel.
Independent self-calibration loops are provided for each of the phase shifters $\phi_{1}$, $\phi_{2}$, $\phi_{3}$, implemented in an all-digital fashion that facilitates fast synthesis. The perturbation frequency, $\omega_{\text{p}}$, is set to a low frequency of 30 rad/s generated using a LUT with 128 entries. Each entry in the LUT is 17-bit long with the first two bits representing sign and magnitude. The HPF cut-off is set as 5 rad/s, to capture the slow variations in the $BF_{\text{out}}$. The accumulator gain of each loop is set as follows: $a_{\phi\text{1}}=15$ , $a_{\phi\text{2}}=20$, and $a_{\phi\text{3}}=25$ to achieve faster convergence. Each loop provides a 16-bit digital output representing the phase code word for the corresponding phase shift control. The $BF_{\text{out}}$ captured from BFIC is normalized to a 16-bit input ranging from -31 to +32, where MSB bit represents the sign, next 5-bits represent the integer value, and lower 10-bit for the fractional part. The calibration loop occupies 0.026 $\text{mm}^{2}$ area per PS while consuming 1.5 mW.
Figure 5: Measurement setup for deformation correction verification and radiation pattern. a. Illustration of test setup. b. Test bench for over-the-air measurements. c. Calibration loop reduces deformation error to $<1.5^{\circ}$ from $7^{\circ}$; corresponding change in phase shifter codes $\phi_{1}$, $\phi_{2}$, $\phi_{3}$ are captured d. Radiation pattern without deformation and with deformation after correction at $0^{\circ}$ Azimuth. e. (left) weight measurement of $2\times 2$ tile with BFIC and its connections, (right) Measurement setup for worst-case deformation with 38 cm radius of curvature by tiling four $2\times 2$ conformal arrays.
3.3 Measured results
Figure 5a illustrates the test setup employed to validate the BFIC and the proposed dynamic deformation correction technique. The wideband modulated signals are generated in MATLAB and transmitted through a Pasternack PE9887 horn antenna using DAC on Xilinx ZCU216 RFSoC. The 3D-printed array, along with the BFIC, is mounted on the DAMS D6025 antenna measurement platform as shown in Fig. 5b. A 4.2GHz differential clock is provided from ADF4372 EVM (wideband synthesizer) to the internal LO and sampler clock generation circuit of the BFIC. The bias voltages for both the BFIC and the self-calibration loop are provided using a DAC81416 EVM (DAC) and the serial control bits to the on-chip serial to parallel interface (SPI) registers being applied from a Digilent Analog Discovery 2 board. The quadrature $BF_{\text{out}}$ outputs post combining from BFIC is captured on the Xilinx ZCU216 ADCs and subsequently converted to a 16-bit digital word using a mapping function. The digital $BF_{\text{out}}$ is then fed to the calibration loop, interfaced through a Waveshare QFN-64 programmable adapter socket. After each iteration, the new control bits for each phase are captured from the respective calibration loops and updated to the serial control bit stream. This is then fed to the SPI control registers using Digilent Analog Discovery II. The entire data capture loop is automated through the power automation tool [50].
To assess the worst-case bending and verify the effectiveness of the calibration technique, a $4\times 4$ array is deformed to a curvature with a radius of 38cm using a custom-designed antenna holder as shown in Fig. 5e. Within this array, three tiles are disabled, leaving only a $2\times 2$ tile active under deformation, resulting in a maximum beam pointing error of $7^{\circ}$. One antenna is kept as reference (REF) while the calibration loop is connected to the other three antennas to optimize the corresponding PS $\phi_{1}$, $\phi_{2}$, $\phi_{3}$. Fig. 5c shows phase control codes are auto-adjusting to achieve the optimal beamformed output under this deformation. Each loop is settled to minimize the beam pointing error $<1.5^{\circ}$. This error can be further reduced by increasing the PS resolution. The 2dB gain reduction after beam-point error correction is due to the antenna holder and the change in Line-of-Sight (LOS). The phase control codes are properly initialized so that the calibration loop doesn’t get stuck at side-lobes (local maxima). Figure 5d shows the radiation pattern without deformation and with deformation after self-correction. The FoV enhancement after deformation is evident from the radiation pattern.
[…]
Figure 6: a. The $2\times 2$ conformal array tile back side, the receiver IC and its connections, are printed using Cu ink on an AP layer, chip micrograph of beamforming receiver, along with the calibration loop. b. Circular patch antenna printed on Ninjaflex using silver ink. c. Azimuth and d. Elevation patterns of the $2\times 2$ conformal phased array. e. S11 and conversion gain of the single receiver channel.
4 DATA AVAILABILITY
The datasets generated and additional results are available from the corresponding author on reasonable request
5 ACKNOWLEDGEMENTS
This material is based on research sponsored, in part, by Air Force Research Laboratory under agreement number FA8650-20-2-5506, as conducted through the flexible hybrid electronics manufacturing innovation institute, NextFlex, Murdock Foundation, Washington Research Foundation, and WSU Office of Commercialization. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon.
The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government.
Authors acknowledge Dr. Robert Dean at Auburn University for his help with the antenna hold and Sonja Gerard at OEIGraphics for help with Figures 1 and 2.
6 AUTHOR CONTRIBUTIONS
This project was conceived by S.G., S.P., J.N., and J.W. The antenna was designed and additively printed by T.D. and K.K with inputs from P.L.. The Cu ink was developed and verified by A. I. and S.R. The receiver front-end was designed by S.P., A.M., Q.X., E.G., and A.R. with inputs from S.G. and S.S. The discrete-time combiner was designed by E.G., S.P., and A.R. with inputs from S.G.. The phase-interpolator and time-interleaved clock generator was designed by S.P., and Q.X. with inputs from S.S. and S.G.. Calibration loop, design, and validation was done by S.B. and S.P. with inputs from S.G. Measurements of the self-calibrated array were performed by S.P. with inputs from S.G.. S.P., A.R., and S.G. are equal contributors for preparing the manuscript.
7 COMPETING INTERESTS
The authors declare no competing non-financial interests.
8 SUPPLEMENTARY SHEETS
8.1 Environmentally Friendly Alternative for Conventional Subtractive Processes
While lightweight conformal arrays are convenient in many applications, large-scale production of these arrays raises major concerns over chemical waste and water pollution during fabrication. Various steps in the traditional subtractive process result in substantial chemical waste generation, including spent etchants, contaminated rinse water, filter sludges, and discarded chemicals. Spent etchant, a strong acid or base, is used to remove unwanted copper from PCBs. Plating bath wastewater is generated from the electroplating process used to deposit metal layers onto PCBs. Rinse water, used to rinse the PCBs after each fabrication step, contains a variety of acids and bases. Most of these fabrication plants are situated near water sources and require substantial amounts of water for each step. This chemical waste can contaminate soil and water, posing serious environmental challenges.
Amid these concerns, additive printing, also known as 3D printing, is an appropriate solution for printed circuit fabrication, offering numerous advantages over conventional subtractive manufacturing methods. PCB design using additive printing does not require any chemical processing, as the entire circuit board can be printed using an inkjet printer. Additive printing requires only the material needed to create the PCB, significantly reducing waste and eliminating various chemical processes. Furthermore, additive printing allows for the creation of complex, thermally conductive structures within the PCB itself, improving the overall thermal performance of the array. While the initial investment in additive printing can be significant, the long-term cost benefits are substantial. The reduced material waste, fewer processing steps, and shorter development cycles translate to lower production costs over time.
8.2 Single tile $2\times 2$ Receiver Array Performance
Figure 7: Measured beamforming RX performance including a. Beamforming gain over wideband modulated input with 1-element and 4-element b. EVM for 16-QAM modulation c. out-Of-band IIP3 d. in-band IIP3 e.1-dB compression point.
Figure 7a shows the wideband beamforming capabilities of the proposed receiver array. This measurement uses the same test bench shown in Fig. 5b. The wideband signal is generated in MATLAB and transmitted using ZCU216 DAC from the Pasternack PE9887 horn antenna at $0^{\circ}$ angle. An $\approx$12dB beamforming gain is observed when all four elements are enabled. The gain profile has a reduced gain at lower frequencies (for both 1-element and 4-elements) due to the high-pass response of the ZCU216 ADC baluns connected at the BFIC output. The measured constellation plot in Fig. 7b shows that the receiver supports 16-QAM modulation with an error vector magnitude (EVM) of 7.2% while supporting 160 Mbps.
Figures  7c-eshow the linearity performance of a single-channel BFIC. The measured 1-dB compression point is -16 dB, which is in reasonable agreement with the simulation results. The in-band IIP3 shown in Fig. 7d is measured using a two-tone test where the first tone is placed at a frequency offset $\Delta f$ from LO and the second tone at $52\Delta f-$ 5MHz offset from LO such that the IM3 product will fall at 5 MHz. Measured in-band IIP3 is -6 dBm for a gain setting of 7 dB. The out-of-band IIP3 shown in Fig. 7d was also measured using a two-tone test with the first tone at an offset of 250 MHz from the LO and second tone at 370 MHz offset from LO such that the IM3 distortion component will fall at 30 MHz. The measurement shows an increase in out-of-band-IIP3 from in-band-IIP3 as it reaches +4 dBm for a gain setting of 7dB.
The power breakdown for the proposed BFIC is given below, the four RF front-end units consuming highest power. The LNTA, passive mixer, TIA, and buffer together with few test circuits consume  41mW in the total power. Clock drivers, LO Generation, LO PS, sampler clock generation together with the PI and the time-interleaved sample-and-hold clocking consumes 27 mW. The charge-domain summer which includes a 2-stage internally compensated OTA consumes 6 mW including both the quadrature paths. The SPI registers only consume 26 $\mu$W. The 3 variable calibration loop consumes 4.5 mW.
Radiation pattern measurement uses the same test setup with the receiver array and is rotated from $-90^{\circ}$ to$+90^{\circ}$ using the DAMS D6025 (with $5^{\circ}$ resolution) to capture the received power level at each angle. The receiver array is flipped $90^{\circ}$ to capture the radiation pattern in the elevation plane using the same test setup. Figure 6c-d shows the azimuth and the elevation radiation patterns. The test setup limited measurement of the radiation pattern and efficacy deformation induced beam pointing error correction in an arbitrary plane ($El\neq 0^{\circ}$ and $Az\neq 0^{\circ}$).
Appendix A Deformation over a curvature with radius R
Figure 8: Deformation of a uniform linear array over a curvature with radius R.
This section estimates the additional path length required for a uniform linear array (ULA) when deformed over a curvature with radius R, in terms of R. We assume a circle with radius R for the analysis. The transmitter is at an angle $\theta$. Fig. 8a shows a 4-element uniform linear array (ULA) with inter-element distance, d. The ULA deformed over a curvature with radius R is shown in Fig. 8b. Consider the isosceles triangle in Fig. 8c, the additional path length required for the signal to reach second element is $\Delta d$. We can express the chord length C as:
$C=2Rsin(\frac{\phi}{2})$ (1)
where $\phi$ is the angle between two consecutive elements. From Fig. 8(c) the path-length can be expressed in terms of chord length C as:
$\Delta d=Ccos(\Omega)$ (2)
Consider the following relations $\Omega+\Psi+\theta=180$ and $\Omega=90-\theta+\phi/2$ and substituting in (2) yields:
$\begin{split}\Delta d&=2Rsin(\frac{\phi}{2})cos(\Omega)\\ &=2Rsin(\frac{\phi}{2})cos(90-\theta+\frac{\phi}{2})=2Rsin(\frac{\phi}{2})sin(\theta-\frac{\phi}{2})\\ &=Rcos(\theta-\phi)-Rcos(\theta)\end{split}$ (3)
For N-element conformal phased array,
$\Delta d=Rcos(\theta-\phi_{n})-cos(\theta)$ (4)
where $\phi_{n}$ is the angle between successive element.
The array factor of the conformal phased array can be expressed as:
$AF=e^{-j\cdot 2\pi f\cdot R\cdot cos(\theta)}\sum_{i=0}^{N-1}e^{j\cdot 2\pi f\cdot R\cdot cos(\theta-\phi_{n})}$ (5)
Appendix B Gradient Estimation using self-calibration loop
The adaptive update rule for the phase shifter is given below:
$\phi[n]=a_{\phi}sin[\omega_{\mathrm{p}}n]+a_{\phi}A_{v,\phi}\sum_{n=1}^{N}(BF_{\mathrm{out}}[n-1]\cdot sin[\omega_{p}n+\Psi])$ (6)
For simplicity, we consider a ULA. The beamformed output of the ULA after baseband beamforming can be expressed as:
$BF_{\text{out}}=\sum_{i=0}^{N-1}A_{\text{n}}e^{j2\pi(f_{\mathrm{RF}}-f_{\mathrm{LO}})t}\cdot e^{j\phi_{\text{ant,n}}}\cdot e^{j\phi_{\text{n}}}$ (7)
To estimate gradient of $BF_{\text{out}}$ using conventional gradient descent method, we can take the derivative of $BF_{\text{out}}$. This requires the derivative of each channel output, thus necessitating the digitization of each channel using highly linear, power-hungry ADCs.
In the proposed loop, the gradient is estimated using array-level information. The gradient estimation is illustrated below. For simplicity, we are considering a 2-element array with a single calibration loop. The phase-shifter code-word can be expressed as:
$P_{\phi}=\hat{P_{\phi}}+a_{\phi}sin[\omega n]$ (8)
where $\hat{P_{\phi}}$ is the estimated phase code-word after each iteration. Expressing BFout as a function of this phase-shifter code-word as follows:
$BF_{out}=f(P_{\phi})=f(\hat{P_{\phi}}+a_{\phi}sin[\omega n])$ (9)
The gradient can be calculated by considering the Taylor series expansion $f_{a}(x)=f(a)+\partial_{x}f(a)(x-a)$, where $\partial$ is the partial-differential operated on the estimated phase-shifter code-word:
$\begin{split}f_{\hat{P_{\phi}}}(P_{\phi})&=f(\hat{P_{\phi}})+\partial_{P_{\phi}}f(\hat{P_{\phi}})\cdot(\hat{P_{\phi}}+a_{\phi}sin[\omega n]-\hat{P_{\phi}})\\ &=f(\hat{P_{\phi}})+\partial_{P_{\phi}}f(\hat{P_{\phi}})\cdot(a_{\phi}sin[\omega n])\end{split}$ (10)
The first term in (10) is a DC term and the second term is a slow varying sinusoidal. It is straight forward to have a HPF to extract the slow varying term which includes the gradient information.
$HPF_{out}=\partial_{P_{\phi}}f(\hat{P_{\phi}})(A_{HPF}\cdot a_{\phi}\cdot sin[\omega n+\Omega_{HPF}])$ (11)
where $A_{\text{HPF}}$ and $\Omega_{\text{HPF}}$ are the magnitude and phase changes due to HPF. The filter cut-off frequency is chosen such that this term is minimally distorted. The HPF output is now multiplied by the LUT perturbation as follows:
$\begin{split}y_{mul}&=a_{\phi}sin[\omega n]\cdot\partial_{P_{\phi}}f(\hat{P_{\phi}})\cdot(A_{HPF}\cdot a_{\phi}sin[\omega n+\Omega_{HPF}])\\ &=a_{\phi}^{2}A_{HPF}\cdot sin[\omega n]\cdot(sin[\omega n+\Omega_{HPF}])\cdot\partial_{P_{\phi}}f(\hat{P_{\phi}})\\ &=X(n)\partial_{P_{\phi}}f(\hat{P_{\phi}})\end{split}$ (12)
where $X(n)$ is the time-varying factor. The multiplier output contains gradient information of $f(\hat{P_{\phi}})$ at $\phi$. The loop estimates gradient from array level information without element level read out. The integrator followed by the multiplier can find the optimum $\phi$ which leads to maximum $f(\hat{P_{\phi}})$